An individual has $50 to spend, and can buy a maximum of 45 litres of gas or 35 bottles of beer.

a. Sketch the budget constraint and choice set.

b. What are the prices of gas and beer?

c. Instead, given the same prices as above, suppose anyone buying gas receives a 10% discount for all purchases over 20 litres. How many bottles of beer can a person buying 20 litres of gas purchase? How much gas can

someone buying no beer purchase? Sketch this budget line with the discount.
Answer #1

a) Income = $50

Maximum gas = 45

Maximum beer = 35

Taking beer on X axis and gas on Y-axis, the intercept of budget constraint on X-axis would be 35 and intercept of Y-axis would be 45.

b) Price of gas = 45 / 50 = 1.11

Price of beer = 35 / 50 = 1.42

c) As there is 10% discount after 20 units of gas. When 20 units of gas are purchased, 19 units of beer can be purchased.

Due to the 10% gas discount, its intercept on Y-axis shifts upwards to 50 units as it makes it price of $1 per unit after discount.